CAT Chat Minutes
July 19, 2004
General Information
Machine Stats
Delivered beam: 1127
Percentage of scheduled time: 98.53
Downtime during period: 16.8 hrs.
Meantime between faults (MTBF): 51.21 hrs.
Faults:
| #31 07/13 16:00 to 7/14 12:37 20.62 | DP-IOC DSP Event | 0.75 Beam lost due to Data Pool IOC problem, refill |
| #32 07/14 13:33 to 07/16 10:56 45.56 | FE EPS trip (33ID) [SI] | 0.50 Human error, trip reset, refill, resume top-up. |
| #33 07/16 11:25 to 07/16 15:39 4.22 | 2-ID RAD Trip [UES] | 0.39 Spurious Radiation Monitor Trip, refill, top-up. |
Beamtime Downtime Summary by APS Groups, 07/12/2004 to 07/19/2004:
DIN |
Date |
Beamline |
Status |
Duration |
Group |
Description |
Resolution |
429 |
07/12/04 |
31-ID |
Open |
20 min |
ASD-VAC |
User reported FE trip, this has been an intermittent problem. Went to SR mezzanine to open BIV after checking vacuum levels. |
Fault reset per standing order. ASD-VAC will investigate causes of trips during next shutdown. |
430 |
07/12/04 |
24-ID |
Open |
2 hr 56 min |
ASD-SI |
C station door will not open. PSS on-call tech troubleshooting the problem. |
|
431 |
07/13/04 |
18-ID |
Closed |
20 min |
ASD-VAC |
BP ( Elena ) from 18-ID reported #67 fault PS2_NOSWITCH-S. Followed Standing Orders & reset fault - it tripped again with the same fault. Reset Serious fault again. |
Reset faults per Standing Order. Relayed information to responsible group. |
432 |
07/14/04 |
31-ID |
Closed |
1 hr 5 min |
ASD-ME |
FM3 DP fault at 05:51 according to log on FE-EPS touch screen on mezzanine. |
Fault reset per standing order. ASD-ME will investigate causes of dP trips during next shutdown. |
433 |
07/14/04 |
18-ID |
Open |
21 min |
ASD-VAC |
Fault 67 PS2_NOSWITCH-S. |
Repairs are scheduled for 7/19/04. Work Request 15383 submitted and approved. |
436 |
07/16/04 |
33-ID |
Open |
5 hr 30 min |
ASD-SI |
FEEPS trip due to bad Love controller FM2-DP. |
Replace faulty LOVE controller. |
437 |
07/17/04 |
16-ID |
Open |
4 hr 51 min |
Serious Fault #112 (P8B_MS1_NOSWITCH) occurred. I have reset the fault three times (per FC standing order). I will contact Jon Hawkins (PSS on-call engineer). |
ASD-SI adjusted air cushion stops to allow proper travel of shutter cylinder. |
|
437 |
07/17/04 |
16-ID |
Open |
34 min |
ASD-SI |
Serious Fault #112 (P8B_MS1_NOSWITCH) occurred. I have reset the fault three times (per FC standing order). I will contact Jon Hawkins (PSS on-call engineer). |
ASD-SI adjusted air cushion stops to allow proper travel of shutter cylinder. |
Machine Studies Schedule for Run02-09, 2004, July 19th, 0800 - July 21st, 0800:
Time |
Description |
Studiers |
SRStatus |
Monday, 7/19/04 |
|||
0800-0805 |
Record XBPM data for user orbit with TopUp running |
OPS |
SR Beam |
0805-0900 |
Waveguide switch from RF3/RF4 to RF2/RF4 |
RF Group |
No Beam |
0805-1000 |
Upgrade power supply for MPEOB strobe in Zone C |
McNamara |
No Beam |
0805-1000 |
Swap out corrector power supplies in Zone C for reg upgrade |
McNamara |
No Beam |
0805-1000 |
Install and test a fix for the A:H3 corrector problem |
Lenkszus |
No Beam |
0805-0900 |
Reboot iocs40fb with a database that supports automated reflective memory testing |
Lenkszus |
No Beam |
0805-1000 |
Test DACI software with SR PSCU |
Hillman |
No Beam |
0805-1000 |
Test the K1 relay a potential external interlock problem with the Raws supplies starting in sector 6:7 and 36:37 |
Sprau |
No Beam |
0900-1200 |
Wireless testing Bldg. 420 |
Stejskal |
No Impact? |
0805-0900 |
Install TLD badges in the LEUTL tunnel on the beam pipe |
Vacca |
No impact |
0805-0900 |
Replace Gespac ADC board for the RF gun heater supply |
Hillman |
No Inj |
0805-0930 |
Reboot linac iocs : ioclirfgun , ioclic1, ioclic4 to add pvs to monitor PS setpoint changes relative to conditioning |
Varotto |
No Inj |
0830-1030 |
Test cables and replace filter comparators at 10BP2, 15BP2, and 15BP3 |
Erwin |
Access Zone B |
0830-0930 |
Reboot iocs3VP to incorporate vacuum pump database additions |
Quock |
No Beam |
0830-1030 |
Terminate existing ACIS cable and power cable for a future radiation monitor at the 26-ID Ratchet Door Column |
Vacca , Forrestal |
No Beam |
0830-0930 |
Reboot iocfe31 |
M. Smith |
No Beam |
0830-0930 |
Replace red beacon at L1:MO:PF |
Cours |
No Inj |
0830-0930 |
Replace rf source backup generator |
Cherbak |
No Beam |
0830-1230 |
Replace the BOOSTER TO LEUTL radiation stop solenoid air manifold, verify the wiring of the ACIS limit switches, perform a `mini' validation of the affected ACIS |
Markovich , Gagliano |
Access Booster (AA mode) |
0830-1130 |
Replace burnt out magnet strobes lights |
Hillman |
Access Booster |
0830-1000 |
Bob Keithley to obtain measurements of BTX beamline for future ODR test station |
Lumpkin |
Access Booster |
0845-1000 |
Check operation of PS2 at 18ID front end. Requires PSS validation of switches on PS2 |
Gagliano |
Access Zone C (AA mode) |
0930-1130 |
Operator training - PAR(0700-1500) |
Glenn , Jones |
No Inj |
0900-1000 |
Replace rotary encoder on upstream end of the 3IDds insertion device |
Merritt |
Access Zone A |
0900-1000 |
Reboot siocops (soft ioc ) to load sequence program to automate testing of reflective memory. |
Varotto |
No Impact |
0900-1000 |
Add bus analyzer board to sector 5 (iocs5vp) to collect data |
Varotto |
No Impact |
0900-0930 |
Remove the VTM200 10/100Mb media converter undergoing field testing at iocs1bpm and restore connections to the Cabletron Ethernet media converter |
Laird |
No Beam |
1000-1130 |
Test cable and replace filter comparator at 35BP2 |
Erwin |
Access Zone E |
1000-1030 |
Install cable in iocmtime racks to bring P0 to timing patch panel |
Lenkszus |
No Beam |
1230-1300 |
Beam recovery. |
OPS |
Stored beam and injection |
1300-1500 |
Reproduce Nan problem with control law. |
Emery Lenkszus |
Stored beam nad injection |
1500-1600 |
CPU mode testing. |
Smith Emery |
Stored beam and injection |
1600-1700 |
Test servo motor. |
Stein Xu Sajaev |
Stored beam and injection |
1700-2100 |
Collect nbbpm spectrum. Parallel to RF phase... |
Sun |
Stored beam, no injection |
1700-2100 |
RF phase anf timing jitter measurements in linac . |
Borland |
No injection |
2100-2400 |
BPM tests. |
Singh |
Stored beam and injection |
Tuesday, 7/20/04 |
|||
0000-0600 |
Canted undulator xbpm studies. |
Decker |
Stored beam and injection |
0600-0900 |
Prepare for 1296 bunches test. |
Sereno Harkay Sajaev Decker et al. |
Stored beam and injection |
0900-1700 |
1296 bunch mode test with shutters open. Floor coordinator support required. |
Randall |
Stored beam and injection |
1000-1200 |
Operator Training - Linac (0700-1500) can be concurrent to the User portion of 1296 bunch operation |
Lesson, Jones |
No injection |
1700-2100 |
Bunch purity study. |
Sereno Borland |
Stored beam and injection |
2100-2400 |
Bunch purity vs. number of bunches study. |
Sereno |
Stored beam and injection |
Tuesday, 7/21/2004 |
|||
0000-0700 |
Xbpm alignment work. |
Decker |
Stored beam and injection |
0100-0300 |
Operator Training - Booster. Parallel to xbpm alignment. |
Forth |
Limited Inj |
0700-0800 |
Ready for 0+24X1 fill pattern, low emittance , TopUp |
OPS |
Stored beam and injection |
Kevin Randall discussed the upcoming bunch mode test on July 20th. User asked if problems will recur. Expected lifetime will be >100 hours. Bunch spacing is 2.85 nanoseconds. Detailed schedule is as follows:
Time |
Status |
0600 - 0900 |
Machine Prep for 1296 bunch test: Sereno, Harkay, Sajaev, Decker, et. al. |
0900 - 1700 |
1296 bunch mode test with shutters open: floor coordinator support required. |
Annual Report & Activity Report - R. Fenner : Discussed the APS activity reports.
Safety E. Chang: A laser eye injury has occurred at LANL. Suggests the use of a beam on indicator to prevent occurrences. Stressed adherence to procedures. A user asked how someone would mistake a laser for being off when it is in fact on. Elroy Chang stated that when in doubt, user should always assume laser is on, and never look into the beam. Elroy also suggested turning down laser intensity so that less optically-dense goggles can be used to locate the beam on a fluorescent card.
In an emergency it is preferred that you call 911. This makes it easier for the laboratory's emergency personnel to locate you. From a cell phone you can also call 630-252-1911. This is not the preferred method of contact, since it is more difficult for emergency personnel to determine your location.
Beware of counterfeit valves from Hunt Valves, Inc. produced between 1992 and 2001. A user asked if problems were expected with the valves? An investigation is underway.
Richard Hislop gave an update on CNM construction. A portion of Kearney Road is fenced off between Rock Road and Bluff Road . The gates are still open. The gates will be closed when construction is completed on Bluff Road to the east of the APS. There will be no traffic through the parking lot (431Z area).